OpenStudy (anonymous):

Simplify, (-3ab^3)^4 i'm not sure about this one.. help please. :(

OpenStudy (anonymous):

Apply 4 to every term..

OpenStudy (anonymous):

\[(-3)^4 (a)^4 (b^3)^4\]

OpenStudy (anonymous):

-12a4b7?

OpenStudy (anonymous):

\((-3)^4\) = ???

OpenStudy (anonymous):

oh..-81a4b7?

OpenStudy (anonymous):

\[(-3)^4 = -3 \times -3 \times -3 \times -3\]

OpenStudy (anonymous):

oh. positive 81!

OpenStudy (anonymous):

81a4b12?

OpenStudy (anonymous):

now check for b power ..

OpenStudy (anonymous):

Yes you are right...

OpenStudy (anonymous):

thankyou!

OpenStudy (anonymous):

Welcome dear..